a geostationary sattelite is orbiting the earth at the height 6R above the surface of earth, R being the radius of earth. the time period of another satellite at a height of 2.5R from the surface of the earth is,
A)10h  B)6/root2h  C)6h  D)6root2h

Grade:10

1 Answers

Arun
6 years ago
Distance of satellite from center of earth are 7R and 3.5 R respectively
Time period of geostationary satellite is 24 hours
(T2/T1) =(R2/R1)3/2
T2 = 24 (3.5R/7R)3/2
T2 = 6\sqrt2 hours